# Latent space models for grouped multiplex networks

### v1

@@ Abstract
Complex multilayer network datasets have become ubiquitous in various applications, such as neuroscience, social sciences, economics, and genetics. Notable examples include brain connectivity networks collected across multiple patients or trade networks between countries collected across multiple goods. Existing statistical approaches to such data typically focus on modeling the structure shared by all networks; some go further by accounting for individual, layer-specific variation. However, real-world multilayer networks often exhibit additional patterns shared only within certain subsets of layers, which can represent treatment and control groups, or patients grouped by a specific trait. Identifying these group-level structures can uncover systematic differences between groups of networks and influence many downstream tasks, such as testing and low-dimensional visualization. To address this gap in existing research, we introduce the GroupMultiNeSS model, which generalizes the previously proposed MultiNeSS model of multiness. This model enables the simultaneous extraction of shared, group-specific, and individual latent structures from a sample of multiplex networks — multiple, heterogeneous networks observed on a shared node set. For this model, we establish identifiability, develop a fitting procedure using convex optimization in combination with a nuclear norm penalty, and prove a guarantee of recovery for the latent positions as long as there is sufficient separation between the shared, group-specific, and individual latent subspaces. We compare the model with MultiNeSS and other models for multiplex networks in various synthetic scenarios and observe an apparent improvement in the modeling accuracy when the group component is accounted for. Experiment with the Parkinson’s disease brain connectivity dataset of TaoWu2017 demonstrates the superiority of GroupMultiNeSS in highlighting node-level insights on biological differences between the treatment and control patient groups.
@@ 1 Introduction
Complex multilayer network datasets have become ubiquitous in many statistical applications, with neuroscience, genetics, social sciences, and economics, among others. Notable examples of such datasets include brain connectivity networks observed in a group of patients, protein-protein interaction networks observed across multiple tissues, or trade networks between countries for various goods. The development of statistical methodologies for modeling such data has been an active area of research in recent years. This led to natural generalizations of classical single-layer models, such as the stochastic block model (SBM) (Holland1983), the random dot product graph (RDPG) (Young2007; Athreya2017), and the unifying model of MaMaUnifModel2020 to their respective multilayer versions (MSBM2015; Jones2021MRDPG; xuefei2020flexiblelatspace). Other lines of work focused on Bayesian latent space approaches (Gollini2016; SalterTownshend2017; DAnegelo2019; Sosa2021), and some attempted to develop novel frequentist approaches, such as the Common Subspace Independent Edge Model (COSIE) (Arroyo2019) and the MultiNeSS model (multiness; TianYinqiu2024). Despite their methodological differences, all of these models address the same key question of how to model the common structure shared by all layers, while incorporating individual network-specific information. For example, multilayer SBM does that by assuming the community labels of the nodes to be the same across layers but allowing the community connectivity probabilities to differ, COSIE assumes all expected layer adjacency matrices lie in a common subspace but can differ within it, and MultiNeSS allows for decomposition of each layer’s latent space into a disjoint union of the shared and individual subspaces.
In this work, we consider the natural follow-up question of how to account for the additional structures specific only to subgroups of network layers within the collection and how to separate these group-level information pieces from the information shared by all the layers. For us, this question is mainly motivated by neuroscience studies, where it is of interest to identify brain regions that exhibit abnormal connectivity patterns in patients with a given disease, such as ADHD (DiffADHD2017), Alzheimer (DiffAlzheimer2024), or Parkinson’s disease (TaoWu2017), compared to brain networks of healthy controls. These works, as well as other neuroscience studies, mainly base their findings on permutation tests applied to appropriate descriptive statistics of network groups, such as average connectivity or average path length. This approach has an apparent limitation, as descriptive statistics are often insufficiently expressive and cannot account for all the complex underlying connectivity patterns in the network, thus providing tests of very low power. On the other hand, most existing statistical works addressing this problem assume a latent space network-generating model and develop hypothesis tests to assess whether the latent positions of nodes in the two samples are drawn from the same distribution, usually up to an orthogonal rotation or scaling (TangSemiparamNetworkTest; nguen2024networktwosampletestblock; macdonald2024mesoscale). Although these approaches help determine whether there exists a statistically significant difference in the two network samples, they are incapable of providing quantitative or visual node-level insights on where this difference comes from – the information that could be crucial for many downstream analyses, such as the identification of malfunctioning brain regions or connections between them.
@@
To address this question, we propose a latent space model for a collection of multiplex networks exhibiting group-level latent structures. It generalizes the previously proposed MultiNeSS model by allowing the latent positions of nodes within each layer to comprise not only a component shared across all networks and an individual component specific only to this layer, but also an additional group component shared by the networks only within its group. We develop a computationally efficient fitting procedure for the proposed model, building upon the estimation techniques for MultiNeSS, and establish theoretical guarantees for it under the Gaussian edge-entry model. Through simulation studies and applications to real-world data, we show that incorporating the group component significantly improves modeling accuracy. Our results highlight the importance of accommodating group-level structure in latent space models and provide a versatile tool for analyzing grouped network data.
@@
The rest of this manuscript is organized as follows. In Section 2, we formally describe the proposed model and state the identifiability conditions for its parameters. In Section 3, we formulate the fitting procedure and propose a cross-validation approach to tuning the algorithm’s hyperparameters. In Section 4, we establish consistency for the fitting algorithm under the Gaussian edge entry distribution. In Section 5, we provide numerical simulations on synthetic data to justify the efficacy of our method, and in Section 6, we apply it to the real-world dataset in TaoWu2017 and study brain connectivity differences in patients diagnosed with Parkinson’s disease. Finally, in Section 7, we discuss the future developments of our approach.
2 Model for grouped multiplex networks
2.1 GroupMultiNeSS
@@
Here we propose a new latent space model for GROUPed MULTIplex NEtworks with Shared Structure (GroupMultiNeSS). We start by fixing the notation. Consider $M$ undirected networks on a shared set of $n$ nodes and assume that the networks are separated into $K$ groups of sizes $m_{k},k=1,\ldots,K$ so that $M=\sum_{k=1}^{K}m_{k}$. Each of these networks is encoded via a symmetric and possibly weighted adjacency matrix
@@
and we will refer to an edge between nodes $i$ and $j$ in layer $\ell$ of group $k$ as $A_{k\ell,ij}$. A node $i$ within each layer $A_{k\ell}$ is associated with a fixed latent position $x_{k\ell}^{(i)}\in\mathcal{X}_{k\ell}\subset\mathbb{R}^{D_{k\ell}}$ where the latent space $\mathcal{X}_{k\ell}$ is equipped with a symmetric function $\kappa_{k\ell}:\mathcal{X}_{k\ell}\times\mathcal{X}_{k\ell}\rightarrow\mathbb{R}$ capturing similarity between the two input latent positions. Stacked latent positions of all the nodes in $A_{k\ell}$ will be denoted as a matrix $X_{k\ell}\in\mathbb{R}^{n\times D_{k\ell}}$. We assume that conditional on $X_{k\ell}$, the elements of $A_{k\ell}$ are independent and follow an edge entry distribution
$A_{k\ell,ij}\stackrel{{\scriptstyle\text{ind}}}{{\sim}}f(\cdot;\kappa_{k\ell}(x_{k\ell}^{(i)},x_{k\ell}^{(j)}),\phi_{k\ell}),\quad 1\leq i\leq j\leq n,\quad(k,\ell)\in\mathcal{I}.$ (1)
Here, $f(\cdot;\theta,\phi)$ depends on some possible layer-dependent nuisance parameters $\phi$ and a scalar parameter $\theta$ capturing the effect of latent similarity between nodes $i$ and $j$ on the edge connecting them in the corresponding layer. For example, if each $\kappa_{k\ell}$ is the standard Euclidean inner product on $\mathcal{X}_{k\ell}$ such that $x^{\top}y\in[0,1]$ for any $x,y\in\mathcal{X}_{k\ell}$, and $f(\cdot;\theta)=\operatorname{Bernoulli}(\theta)$, we get the Multilayer RDPG model (Jones2021MRDPG). For most applications, it would usually be sufficient to restrict the edge-entry distribution to belong to a one-parameter canonical exponential family
@@
with a natural parameter $\theta$ and a log-partition function $\nu$ as it includes Bernoulli, Poisson, Gaussian, and other distributions capable of modeling nearly all possible edge entry types, including binary, count, and continuous. For now, we do not make any assumptions on the distribution family and state the model for an arbitrary edge-entry distribution for completeness.
Notice that (1) assumes the distribution of layer $A_{k\ell}$ depends on the latent positions $X_{k\ell}$ only through the Gram matrix $\Theta_{k\ell}\in\mathbb{R}^{n\times n}$ with $\Theta_{k\ell,ij}=\kappa_{k\ell}(x_{kl}^{(i)},x_{k\ell}^{(j)})$. In particular, when $\kappa_{k\ell}$ is the standard Euclidean inner product, $\Theta_{k\ell}=X_{k\ell}X_{k\ell}^{\top}$ obtains the recognizable low rank matrix form that widely appears in the latent space modeling literature following the seminal work of (Hoff2002). To emphasize this observation, we rewrite (1) in the matrix form for convenience as follows
@@
The assumption that the layer distribution depends on the latent positions $X_{k\ell}$ only through their Gram matrix $\Theta_{k\ell}$ naturally leads to identifiability issues which will be discussed in section 2.2 in more detail.
@@
Until now, we have treated $X_{k\ell}$ as independent variables and have not imposed any relationship between the latent positions of the nodes between layers. The main assumption of the proposed GroupMultiNeSS model (see Figure 1) is that the latent positions $X_{k\ell}\in\mathbb{R}^{n\times D_{k\ell}}$ contain some common structure $V\in\mathbb{R}^{n\times d_{0}}$ shared by all the layers, a group-specific structure $W_{k}\in\mathbb{R}^{n\times d_{k}}$ shared by the layers only within a given group $k=1,\ldots,K$, and finally, some individual structure $U_{k\ell}\in\mathbb{R}^{n\times d_{k\ell}}$ specific only to a given layer $\ell=1,\ldots,m_{k}$ within group $k$:
$X_{k\ell}=[V\ W_{k}\ U_{k\ell}],\quad(k,\ell)\in\mathcal{I},$ (4)
so that $D_{k\ell}=d_{0}+d_{k}+d_{k\ell}$. Without the group components, that is, if $d_{k}=0$ for each $k=1,\ldots,K$, this model would transform into the previously proposed MultiNeSS model of multiness.
To allow $\kappa_{k\ell}$ capture richer dependencies between the nodes’ latent positions, we assume it as the generalized inner product that takes vectors $x$ and $y$ in $\mathbb{R}^{p+q}$ as input and outputs
@@
where $I_{p,q}$ is a $(p+q)\times(p+q)$ diagonal matrix with first $p$ diagonal values equal to $1$ and remaining $q$ equal to 1-1. In this decomposition, the first $p$ and last $q$ latent dimensions are referred to as assortative and disassortative, respectively. Intuitively, similarity along an assortative dimension increases the edge weight between two nodes, whereas similarity along a disassortative dimension decreases it. For full generality, we also allow each latent component to contain a different number of assortative and disassortative dimensions. That is, we assume $\Theta_{k\ell}$ admits the decomposition
@@ 2.2 Identifiability
Here, we formulate a sufficient condition for the identifiability of the GroupMultiNeSS parameters. For this section only, we assume a slightly more general form of the similarity function $\kappa$ than we did in (5). That is, we assume $\kappa(x,y)=\psi(x^{\top}I_{p,q}y)$ is an invertible scalar function of the generalized inner product. In the context of identifiability, this is a natural extension since two edge-entry distributions $f(\cdot;\theta_{1})$ and $f(\cdot;\theta_{2})$ that do not coincide for any $\theta_{1}\neq\theta_{2}$, will also not coincide at $\psi(\theta_{1})\neq\psi(\theta_{2})$ by invertibility.
Assume first we observed just one network layer $A_{k\ell}$ (and thus just one group $k$). Then it would be natural to assume column-wise linear independence of the latent position matrix $X_{k\ell}=[V\ W_{k}\ U_{k\ell}]$ to guarantee identifiability of the latent dimension $D_{k\ell}$, a standard condition in the context of inner-product latent models. Apparently, it is not enough as we need to observe at least two groups to separate $V$ from $W_{k}$ and at least two networks in each group to be able to separate $W_{k}$ from $U_{k\ell}$, implying $K\geq 2$ and $m_{k}\geq 2,k=1,\ldots,K$. Indeed, if we had only one group, mixing the columns of $V$ and $W_{k}$ would not change the Gram matrix in (6) and thus the edge-entry distribution. Similarly, if we observed only one layer $\ell$ in the group $k$, mixing the columns of $U_{k\ell}$ and $W_{k}$ would also preserve the distribution. Unfortunately, these conditions are still insufficient as we can simultaneously put a subset of $W_{k}$’s columns to each $U_{k\ell},\ \ell=1,\ldots,m_{k}$ and a subset of $V$’s columns to each $W_{k},k=1,\ldots,K$ or to each $\{U_{k\ell}\}_{\mathcal{I}}$ without changing the distribution. To avoid this, intuitively, we need $V$ to contain the maximum information shared by all the layers, and $W_{k}$ to have the maximum information shared by the layers in group $k$ but not present in $V$. Finally, even if all this holds, each latent component would be only identifiable up to an indefinite orthogonal transformation in $\mathcal{O}_{p,q}=\{O\in\mathbb{R}^{(p+q)\times(p+q)}:O^{\top}I_{p,q}O=I_{p,q}\}$, where $p$ and $q$ are the numbers of assortative and disassortative dimensions in the corresponding component, respectively. For example, it is easy to notice that the Gram matrix in (6), would not change from substituting $V$ by $VO$ where $O\in\mathcal{O}_{p_{0},q_{0}}$. The following proposition formulates a sufficient condition that gathers all these considerations in one statement. The proof can be found in section B.1 of Appendix.
@@
Assume $\{f(\cdot;\theta,\phi),\ \theta\in\mathbb{R}\}$ is an identifiable parametric family and $\kappa(x,y)=\psi(x^{\top}I_{p,q}y)$ is an invertible function of the generalized inner product. Under the following conditions, parameters of the GroupMultiNeSS model with edge-entry distribution in (1) are identifiable up to indefinite orthogonal transformations:
@@
    For each $k=1,\ldots,K$, there exist layers $(k,s)$ and $(k,t)$ such that the columns of the matrix
    $\left[V\ W_{k}\ U_{ks}\ U_{kt}\right]$
    are linearly independent.
@@
    $V\ W_{k_{1}}\ W_{k_{2}}\ U_{k_{1}\ell_{1}}\ U_{k_{2}\ell_{2}}]$
    are linearly independent.
That is, under these conditions, if the probability distributions induced by two different parameterizations $\left(V,\{W_{k}\}_{k=1}^{K},\{U_{k\ell}\}_{\mathcal{I}}\right)$ and $\left(V^{\prime},\{W_{k}^{\prime}\}_{k=1}^{K},\{U_{k\ell}^{\prime}\}_{\mathcal{I}}\right)$ coincide, then
$V=V^{\prime}O_{0}$
$W_{k}=W_{k}^{\prime}O_{k},$ $k=1,\ldots,K,$
$U_{k\ell}=U_{k\ell}^{\prime}O_{k\ell},$ $(k,\ell)\in\mathcal{I}.$

### v2

@@ Abstract
Complex multilayer network datasets have become ubiquitous in various applications, such as neuroscience, social sciences, economics, and genetics. Notable examples include brain connectivity networks collected across multiple patients or trade networks between countries collected across multiple goods. Existing statistical approaches to such data typically focus on modeling the structure shared by all networks; some go further by accounting for individual, layer-specific variation. However, real-world multilayer networks often exhibit additional patterns shared only within certain subsets of layers, which can represent treatment and control groups, or patients grouped by a specific trait. Identifying these group-level structures can uncover systematic differences between groups of networks and influence many downstream tasks, such as testing and low-dimensional visualization. To address this gap in existing research, we introduce the GroupMultiNeSS model, which generalizes the previously proposed MultiNeSS model of MacDonald et al. (2022). This model enables the simultaneous extraction of shared, group-specific, and individual latent structures from a sample of multiplex networks — multiple, heterogeneous networks observed on a shared node set. For this model, we establish identifiability, develop a fitting procedure using convex optimization in combination with a nuclear norm penalty, and prove a guarantee of recovery for the latent positions as long as there is sufficient separation between the shared, group-specific, and individual latent subspaces. We compare the model with MultiNeSS and other models for multiplex networks in various synthetic scenarios and observe an apparent improvement in the modeling accuracy when the group component is accounted for. Experiment with the Parkinson’s disease brain connectivity dataset of Badea et al. (2017) demonstrates the superiority of GroupMultiNeSS in highlighting node-level insights on biological differences between the treatment and control patient groups.
@@ 1 Introduction
Complex multilayer network datasets have become ubiquitous in many area of applications, including neuroscience, genetics, social sciences, and economics, among others. Examples of such datasets include brain connectivity networks observed in a group of patients, protein-protein interaction networks observed across multiple tissues, or trade networks in multiple commodities between countries; in all these cases, the same set of nodes is shared across multiple networks (layers). A number of statistical methods for such data have been proposed, including natural generalizations of single-layer models, such as the stochastic block model (SBM) (Holland et al., 1983), the random dot product graph (RDPG) (Young and Scheinerman, 2007; Athreya et al., 2017), and the latent variable model of Ma et al. (2020) to their respective multilayer versions (Han et al., 2015; Jones and Rubin-Delanchy, 2021; Zhang et al., 2020). Other lines of work focused on Bayesian approaches to latent space models (Gollini and Murphy, 2016; Salter-Townshend and McCormick, 2017; D’Angelo et al., 2018; Sosa and Betancourt, 2021), and models based on low-rank assumptions, such as the Common Subspace Independent Edge Model (COSIE) (Arroyo et al., 2021) and the MultiNeSS model (MacDonald et al., 2022; Tian et al., 2024). All of these models, in various ways, focus on how to model the common structure shared by all the layers, while accounting for individual layer-specific information. For example, the multilayer SBM assumes the communities to be the same across layers but allows their connection probabilities to differ, COSIE assumes all expected layer adjacency matrices lie in a common subspace, and MultiNeSS decomposes each layer’s latent space into shared and individual subspaces.
While the common structure is the natural starting point for studying multilayer networks, in many applications there are important questions that go beyond common and individual structure. For instance, in the context of brain connectivity, one may want to look at the differences between patients and healthy controls, or between a treatment and a placebo, while separating out both the common structure shared by all humans and the individual variability irrelevant to the scientific question at hand. In this work, we study the question of how to estimate structure specific to groups of network layers within the collection, and how to separate these group-level structures from the structure shared by all the layers.
Typical studies focused on discovering abnormal connectivity patterns in patients with a given disease, such as ADHD (Ghaderi et al., 2017), Alzheimer’s (Wu et al., 2024), or Parkinson’s (Badea et al., 2017), tend to perform permutation tests on descriptive summary statistics of the network groups, such as average connectivity or average path length. These descriptive statistics are often insufficient to account for the full complexity of the networks, resulting in low power of the tests, and are not able to separate out the common structure. Latent space network models have been also used to develop classical two-sample hypothesis tests, testing whether the latent positions of nodes in the two samples are drawn from the same distribution, usually up to an orthogonal rotation or scaling (Tang et al., 2017; Nguen et al., 2024; MacDonald et al., 2024). While these approaches provide valid statistical tests, they are also unable to separate out the common structure before comparing groups, resulting in lower power. Further, both these types of methods do not provide estimates of group structure that could be visualized and compared.
@@
In contrast, here we focus on developing a latent space model specifically for multiplex networks whose layers can be divided into groups. It generalizes the previously proposed MultiNeSS model by allowing the latent positions of nodes within each layer to comprise not only a component shared across all networks and an individual component specific only to this layer, but also an additional group component shared by the networks only within its group. We develop a computationally efficient fitting algorithm for this model, and establish theoretical guarantees under the Gaussian edges model. Through simulation studies and applications to real data, we show that incorporating a group component significantly improves modeling accuracy. Our results highlight the importance of accounting for group-level structure in latent space models and provide a versatile tool for analyzing grouped network data.
@@
The rest of this manuscript is organized as follows. Section 2 presents the model and states the identifiability conditions for its parameters. In Section 3, we describe the fitting procedure and propose a cross-validation approach for selecting hyperparameters. In Section 4, we establish consistency for the fitting algorithm under the Gaussian assumption on edge values. In Section 5, we present numerical simulations on synthetic data to illustrate the efficacy of our method, and in Section 6, we apply it to the Parkinson’s disease brain connectivity data from Badea et al. (2017). Section 7 discusses conclusions and future work.
2 A model for grouped multiplex networks
@@
We start by fixing notation. The observed data are $M$ undirected networks on a shared set of $n$ nodes, separated into $K$ groups of sizes $m_{k}$, $k=1,\ldots,K,$ so that $M=\sum_{k=1}^{K}m_{k}$. Each network is represented by a symmetric and possibly weighted adjacency matrix
@@
where the edge value $A_{k\ell,ij}$ represents the strength of connection between nodes $i$ and $j$ in network $\ell$ of group $k$. Each node $i$, $i=1,\dots,n$ is associated with a layer-specific fixed latent position $x_{k\ell}^{(i)}\in\mathcal{X}_{k\ell}\subset\mathbb{R}^{D_{k\ell}}$, which we stack into a matrix $X_{k\ell}\in\mathbb{R}^{n\times D_{k\ell}}$. We assume that conditional on $X_{k\ell}$, the elements of $A_{k\ell}$ for $i\leq j$ are independently drawn and follow an edge entry distribution
$A_{k\ell,ij}\stackrel{{\scriptstyle\text{ind}}}{{\sim}}f(\cdot;\kappa_{k\ell}(x_{k\ell}^{(i)},x_{k\ell}^{(j)}),\phi_{k\ell}),\quad 1\leq i\leq j\leq n,\quad(k,\ell)\in\mathcal{I},$ (1)
where $\kappa_{k\ell}:\mathcal{X}_{k\ell}\times\mathcal{X}_{k\ell}\rightarrow\mathbb{R}$ is a symmetric function capturing similarity between the two input latent positions, and $\phi_{k\ell}$ is a possible nuisance parameter of the edge distribution $f$. If the modeled networks are free of self-loops, we restrict this assumption to the entries with $i<j$. Many previously proposed multiplex network models have this form; for example, if $\kappa_{k\ell}$ is the inner product, $\mathcal{X}_{k\ell}$ is such that $0\leq x^{\top}y\leq 1$ for all $x,y\in\mathcal{X}_{k\ell}$, and $f$ corresponds to the Bernoulli distribution, we get the Multilayer RDPG model (Jones and Rubin-Delanchy, 2021).
We assume that the edge distribution belongs to a canonical exponential family of the form
@@
with a natural parameter $\theta$ and a log-partition function $\nu$. This family includes Bernoulli, Poisson, Gaussian, and many other distributions, and is suitable for modeling many different edges types across a range of applications, including binary, count, and continuous edge values.
The distribution of edges in layer $A_{k\ell}$ in (1) depends on the latent positions $X_{k\ell}$ only through the Gram matrix $\Theta_{k\ell}\in\mathbb{R}^{n\times n}$ with $\Theta_{k\ell,ij}=\kappa_{k\ell}(x_{kl}^{(i)},x_{k\ell}^{(j)})$; this will naturally raise questions about identifiability, addressed in Section 2.2. In particular, when $\kappa_{k\ell}$ is the Euclidean inner product, $\Theta_{k\ell}=X_{k\ell}X_{k\ell}^{\top}$, the key low rank matrix for many network latent space models starting from the seminal work of (Hoff et al., 2002). To emphasize this natural parameter, we can rewrite (1) in matrix as follows:
@@
2.1 A group latent space structure
Here we propose a new latent space model of the general type (3) which explicitly includes common, group-level, and individual structure. We call it GroupMultiNeSS (GROUPed MULTIplex NEtworks with Shared Structure). The key assumption of GroupMultiNeSS, illustrated in Figure 1, is that the latent positions $X_{k\ell}\in\mathbb{R}^{n\times D_{k\ell}}$ can be decomposed into parts representing common structure $V\in\mathbb{R}^{n\times d_{0}}$ shared by all the layers, group-specific structure $W_{k}\in\mathbb{R}^{n\times d_{k}}$ shared by the layers within a given group $k=1,\ldots,K$, and finally, individual structure $U_{k\ell}\in\mathbb{R}^{n\times d_{k\ell}}$ specific to a given layer $\ell=1,\ldots,m_{k}$ within group $k$; that is,
$X_{k\ell}=[V\ W_{k}\ U_{k\ell}],\quad(k,\ell)\in\mathcal{I},$ (4)
and $D_{k\ell}=d_{0}+d_{k}+d_{k\ell}$. If the number of groups is set to 1, this model recovers the MultiNeSS model of MacDonald et al. (2022) as a special case.
@@ 2.1 A group latent space structure
We set the similarity function $\kappa_{k\ell}$ to be the generalized inner product, defined as
@@
where $I_{p,q}$ is a $(p+q)\times(p+q)$ diagonal matrix with first $p$ diagonal values equal to $1$ and remaining $q$ equal to 1-1. The first $p$ dimensions are referred to as assortative, meaning that higher similarity in this dimension increases the edge weight, and the last $q$ dimensions are disassortative, meaning a higher similarity decreases the edge weight. We allow the number of assostative and disassortative dimensions to be different in each latent component, assuming that $\Theta_{k\ell}$ has the general form
@@ 2.2 Identifiability
In this section, we give a sufficient condition for the identifiability of the GroupMultiNeSS parameters, for a somewhat more general similarity function $\kappa(x,y)=\psi(x^{\top}I_{p,q}y)$, an invertible scalar function of the generalized inner product defined in (5). This is an easy generalization since if $f(\cdot;\theta_{1})\neq f(\cdot;\theta_{2})$ for any $\theta_{1}\neq\theta_{2}$, the distributions will also not coincide whenever $\psi(\theta_{1})\neq\psi(\theta_{2})$.
First consider the case of $M=1$, a single network. The standard identifiability condition for an inner product latent position model is linearly independent columns of the latent position matrix $X_{k\ell}=[V\ W_{k}\ U_{k\ell}]$. However, this is evidently not enough for GroupMultiNeSS, since we need to observe at least two groups to separate $V$ from $W_{k}$ and at least two networks in each group to be able to separate $W_{k}$ from $U_{k\ell}$. We will thus assume $K\geq 2$ and $m_{k}\geq 2$ for all $k$. Further, intuitively we need $V$ to contain all the information shared by all the layers, while $W_{k}$ should contain all of the information shared by the layers in group $k$ but not overlap with $V$, and similarly for $W_{k}$ and $U_{k\ell}$. Finally, as in all inner product models, each latent component is only identifiable up to an indefinite orthogonal rotation, defined by $\mathcal{O}_{p,q}=\{O\in\mathbb{R}^{(p+q)\times(p+q)}:O^{\top}I_{p,q}O=I_{p,q}\}$. The following proposition formalizes this intuition. The proof can be found in section B.1 of Appendix.
@@
Assume $\{f(\cdot;\theta,\phi),\ \theta\in\mathbb{R}\}$ is an identifiable parametric family and $\kappa(x,y)=\psi(x^{\top}I_{p,q}y)$ is an invertible function of the generalized inner product. Assume the following conditions hold:
@@
    For each $k=1,\ldots,K$, there exist $\ell_{1}\neq\ell_{2}$ such that the columns of the matrix $\left[V\ W_{k}\ U_{k\ell_{1}}\ U_{k\ell_{2}}\right]$ are linearly independent.
@@
    $V\ W_{k_{1}}\ W_{k_{2}}\ U_{k_{1}\ell_{1}}\ U_{k_{2}\ell_{2}}]$ are linearly independent.
Then the parameters of the GroupMultiNeSS model (1) are identifiable up to indefinite orthogonal transformations, that is, if the probability distributions induced by two different parameterizations $\left(V,\{W_{k}\}_{k=1}^{K},\{U_{k\ell}\}_{\mathcal{I}}\right)$ and $\left(V^{\prime},\{W_{k}^{\prime}\}_{k=1}^{K},\{U_{k\ell}^{\prime}\}_{\mathcal{I}}\right)$ coincide, then
$V=V^{\prime}O_{0},\ \ W_{k}=W_{k}^{\prime}O_{k},\ \ U_{k\ell}=U_{k\ell}^{\prime}O_{k\ell},$